OpenStudy (anonymous):

]What is the molarity of a 0.25-liter solution containing 52 grams of HBr? 13 M HBr 2.6 M HBr 0.64 M HBr 0.16 M HBr

thomaster (thomaster):

Molarity = mol per liter First calculate how many mol HBr = 52 gram To do this, you need to find the molar mass of hydrogen bromide, and divide 52 with it. Then you'll know how many moles HBr you have If it were 1 liter, you'd had the molarity, but in this case the vollume is 0.25 liter So you have to divide it by 4 to get the molarity.

OpenStudy (anonymous):

thank you so much! Hate online schooling it's ever explained well

thomaster (thomaster):

So did you get the answer?

OpenStudy (anonymous):

yeah it says .38 but that's none of the answers..

thomaster (thomaster):

Let me check

thomaster (thomaster):

What is the molar mass of HBr?

OpenStudy (anonymous):

80.912

thomaster (thomaster):

That's correct. \(\large \dfrac{52}{80.912}=0.643\ mol\) What do you get when you divide that by 4?

OpenStudy (anonymous):

2.6

OpenStudy (anonymous):

thanks(:

thomaster (thomaster):

No you multiplied it.

thomaster (thomaster):

How can 0.643 divided by 4 be 2.6?

OpenStudy (anonymous):

hdjbab wait hold on

OpenStudy (anonymous):

i divided .643 by .25 and it says 2.572 so..

OpenStudy (anonymous):

that's one of the answers so should I go with that?

thomaster (thomaster):

The molarity is in mol per liter since you have 1/4th part of a liter (0.25) you should divide it by 4, not 0.25 because 1/4 = 0.25

thomaster (thomaster):

\(\large\sf\dfrac{0.643}{4}=?\)

OpenStudy (anonymous):

.16

OpenStudy (anonymous):

my bad

thomaster (thomaster):

Haha no problem, you only learn from these little mistakes :)

OpenStudy (anonymous):

lol yup! thanks I'm probably gonna need help with another problem in a sec this is pretty difficult

thomaster (thomaster):

Sure i can help you with any of these kinda questions :)

OpenStudy (anonymous):

A sample of gas has a pressure of 3.2 atmospheres at a temperature of 301 Kelvin. What would be the pressure of the same sample of gas when the temperature is decreased to 285 Kelvin? Show all of the work used to solve this problem.

OpenStudy (anonymous):

lol like this one.

thomaster (thomaster):

It's best to close this question and start a new one with that question :)

OpenStudy (anonymous):

oh ok

thomaster (thomaster):

otherwise the post can get really long and confusing